*100 sets of identical twins and 100 sets of fraternal twins are studied for rates of asthma. Results are: Identical Twins: Both have disease: 4, Neither have it: 61, One has it/other does not have it 35. Fraternal Twins: Both have disease 15, Neither have it: 75, One has it/other does not have it 10. Calculate the Concordance and Disconcordance Rate for the Identical Twins and for the Fraternal Twins.
Answer #1

For Identical twins,

Number of Concordance cases = 4 + 61 = 65

Concordance rate

Number of Disconcordance Cases = 35

Disconcordance Rate

For Fraternal twins,

Number of Concordance cases = 15 + 75 = 90

Concordance rate

Number of Disconcordance Cases = 10

Disconcordance Rate